# A rigidity theorem for complex Kleinian groups

### v1

@@ 5. Rigidity
for all $v_{1},\dots,v_{d-1}\in\operatorname{\mathbb{C}}^{2}$ and all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. By Lemma 4.3, $\Gamma^{\prime\prime}:=\tau^{-1}(\Gamma^{\prime})$ is a $\operatorname{\mathsf{P}}_{\alpha_{1}}$-Anosov (i.e. convex cocompact) subgroup of $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ whose limit set if all of $\mathbb{CP}^{1}$, and hence is a lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. It is straightforward to verify that for all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ and $i\in\{1,\dots,d-1\}$,
$\alpha_{i}(\kappa(\tau(g)))=\alpha_{1}(\kappa(g)),$
where $\kappa$ on the left and the right are the Cartan projections for $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ and $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ respectively. It now follows from Theorem 2.2 that $\Gamma^{\prime}$, and hence $\Gamma$, is a $\operatorname{\mathsf{P}}_{\Delta}$-Anosov subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$.
@@
from $\operatorname{\mathsf{G}}$ to the biholomorphism group of $\Lambda_{1}(\Gamma)$. Since $\Lambda_{1}(\Gamma)$ contains $(d+1)$-points in general position, the only element of $\operatorname{\mathsf{G}}$ acting trivially on $\Lambda_{1}(\Gamma)$ is the identity. Hence $\ker q_{1}$ is trivial. At the same time, since $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})\cong\mathsf{Aut}_{hol}(\Lambda_{1}(\Gamma))$, $q_{1}|_{\operatorname{\mathsf{G}}^{0}}$ is an isomorphism. Thus $\operatorname{\mathsf{G}}=\operatorname{\mathsf{G}}^{0}$. ∎
It now remains to prove Lemma 5.2. To do so, we first observe from the following result of Pozzetti–Sambarino–Wienhard [30] (see also Zhang–Zimmer [37] for the case when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$) that $\Lambda_{k}(\Gamma)$ is a differentiable submanifold.
@@
([30, Prop. 7.1]) Suppose that $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (d1)(d-1)-hyperconvex subgroup. If $\{(z_{n},w_{n})\}$ is a sequence of distinct points in $\Lambda_{\theta_{1}}(\Gamma)$ that converges to $(x,x)$, then
@@
Notice that for every $x\in\Lambda_{\theta_{1}}(\Gamma)$, the set of points in $\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that lie in $x^{2}$ is a projective subspace of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that we also denote by $x^{2}$. Furthermore, for any affine chart $U\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that contains $x^{1}$, $x^{2}\cap U$ is an affine subspace of $U$ that contains $x^{1}$, so we may view $x^{2}$ as a $\operatorname{\mathbb{K}}$-linear subspace of $T_{x^{1}}\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})\cong T_{x^{1}}U$. Thus, in the case that $\operatorname{\mathbb{K}}=\mathbb{C}$ and $\partial\Gamma$ is a two-sphere, Theorem 5.3 immediately implies the following consequence. (For comparison, see [30, Prop. 7.4] and [37, Thm. 1.1].)
@@
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (d1)(d-1)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{1}(\Gamma)$ is a $\mathcal{C}^{1}$-submanifold of $\mathbb{P}(\mathbb{C}^{d})$. Furthermore, for every $x^{1}\in\Lambda_{1}(\Gamma)$, and the tangent space $T_{x^{1}}\Lambda_{1}(\Gamma)\subset T_{x^{1}}\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ is a complex linear subspace.
To further upgrade $\Lambda_{1}(\Gamma)$ to a complex submanifold of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$, we use the following general statement.
@@
By Proposition 2.6, it suffices to prove the lemma for the case when $k=1$. The case when $k=1$ is immediate from Corollary 5.4 and Lemma 5.5. ∎
@@
We will now give a sketch of our original, more dynamical proof prior to the simplification suggested by Andres Sambarino.
@@
By Theorem 3.1, the $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$ is Ahlfors regular and the $\alpha_{1}$-Patterson–Sullivan measure of $\hat{\Gamma}$ on $\mathbb{CP}^{1}$ is mutually absolutely continuous with the Lebesgue measure. Since quasi-conformal maps preserve the Lebesgue measure class, we can then conclude that their pull-backs to $\partial\Gamma$ are mutually absolutely continuous. Then a result of Blayac–Canary–Zhu–Zimmer [5, Thm. 13.2] (see also Sambarino [32, Cor. 3.3.4]) implies that
@@
We then consider the identity component $\mathsf{G}^{0}$ of the real Zariski closure $\mathsf{G}$ of $\Gamma$. We know, by Theorem 4.1, that $\mathsf{G}^{0}$ is simple. We may now use an argument of Dal’bo–Kim [15] to show that $\mathsf{G}^{0}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$. Once we have done so, we can complete the proof of Theorem 5.1 just as above.
@@ Appendix A Reducing to the projective case
Observe then that the smooth embedding $E_{k}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\hookrightarrow\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})=\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{\bar{d}})$ defined by
$E_{k}(\operatorname{Span}(v_{1},\dots,v_{k}))=[v_{1}\wedge\cdots\wedge v_{k}]$
is $\eta_{k}$-equivariant, and is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$.
@@
First, we prove (1). By (VII), the fact that $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-divergent implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-divergent, and $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))=E_{\theta_{k}}(\Lambda_{\theta_{k}}(\Gamma))$. In particular, $\Lambda_{1}(\eta(\Gamma))=E_{k}(\Lambda_{k}(\Gamma))$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-transverse, (III) implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-transverse. Since $E_{\theta_{k}}$ is an $\eta$-equivariant embedding, the fact that every point in $\Lambda_{\theta_{k}}(\Gamma)$ is a conical point for the $\Gamma$-action implies that every point in $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))$ is a conical point for the $\tau(\Gamma)$-action. Hence, $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-Anosov.

### v2

@@ 4. Rigidity
for all $v_{1},\dots,v_{d-1}\in\operatorname{\mathbb{C}}^{2}$ and all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. The standard (Hermitian) inner product $\langle\cdot,\cdot\rangle$ on $\operatorname{\mathbb{C}}^{2}$ induces an inner product $(\cdot,\cdot)$ on ${\rm Sym}^{d-1}(\operatorname{\mathbb{C}}^{2})$ with the defining property
$(v_{1}\odot\dots\odot v_{d-1},u_{1}\odot\dots\odot u_{d-1}):=\sum_{\sigma\in S_{d-1}}\prod_{i=1}^{d-1}\langle v_{i},u_{\sigma(i)}\rangle,$
where $S_{d-1}$ is the symmetric group on the set $\{1,\dots,d-1\}$. It is straightforward to verify that for all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ and $j\in\{1,\dots,d-1\}$,
$\alpha_{j}(\kappa(\tau(g)))=\alpha_{1}(\kappa(g)),$
where $\kappa$ on the left and the right are the Cartan projections for $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ and $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ with respect to $(\cdot,\cdot)$ and $\langle\cdot,\cdot\rangle$ respectively. In particular,
$\alpha_{j}(\kappa(\gamma))=\alpha_{k}(\kappa(\gamma))$
for all $\gamma\in\Gamma^{\prime}$ and $j,k\in\{1,\dots,d-1\}$. It now follows from Theorem 2.2 that $\Gamma^{\prime}$, and hence $\Gamma$, is a $\operatorname{\mathsf{P}}_{\Delta}$-Anosov subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$.
@@
from $\operatorname{\mathsf{G}}$ to the biholomorphism group of $\Lambda_{1}(\Gamma)$. Since $\Lambda_{1}(\Gamma)$ contains $(d+1)$-points in general position (it is the image of the Veronese embedding from $\mathbb{CP}^{1}$ to $\mathbb{CP}^{d-1}$), the only element of $\operatorname{\mathsf{G}}$ acting trivially on $\Lambda_{1}(\Gamma)$ is the identity. Hence $\ker q_{1}$ is trivial. At the same time, since $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})\cong\mathsf{Aut}_{hol}(\Lambda_{1}(\Gamma))$, $q_{1}|_{\operatorname{\mathsf{G}}^{0}}$ is an isomorphism. Thus $\operatorname{\mathsf{G}}=\operatorname{\mathsf{G}}^{0}$. ∎
It now remains to prove Lemma 4.2. To do so, we first observe from the following result of Pozzetti–Sambarino–Wienhard [27] (see also Zhang–Zimmer [35] for the case when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$) that $\Lambda_{1,2}(\Gamma)$ is a differentiable submanifold when $\Gamma$ is (d1)(d-1)-hyperconvex.
@@
([27, Prop. 7.1]) Suppose that $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (d1)(d-1)-hyperconvex subgroup. If $\{(z_{n},w_{n})\}$ is a sequence of distinct points in $\Lambda_{\theta_{1}}(\Gamma)$ that converges to $(x,x)$, then
@@
Thus, in the case that $\operatorname{\mathbb{K}}=\mathbb{C}$ and $\partial\Gamma$ is a two-sphere, Theorem 4.4 immediately implies the following consequence. (For comparison, see [27, Prop. 7.4] and [35, Thm. 1.1].)
@@
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (d1)(d-1)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{1}(\Gamma)$ is a $\mathcal{C}^{1}$-submanifold of $\mathbb{P}(\mathbb{C}^{d})$. Furthermore, for every $x\in\Lambda_{1,2}(\Gamma)$, we have
$T_{x^{1}}\Lambda_{1}(\Gamma)=T_{x^{1}}\operatorname{\mathbb{P}}(x^{2}).$ (5)
In the above corollary, $T_{x^{1}}\operatorname{\mathbb{P}}(x^{2})$ is a complex linear subspace of $T_{x^{1}}\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$. So to further upgrade $\Lambda_{1}(\Gamma)$ to a complex submanifold of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$, we use the following general statement.
@@
By Proposition 2.6, it suffices to prove the lemma for the case when $k=1$. Suppose $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is a (d1)(d-1)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere. By Corollary 4.5 and Lemma 4.6, $\Lambda_{1}(\Gamma)$ is a complex submanifold of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$.
Let $\operatorname{\mathrm{Gr}}_{1}(T\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d}))$ denote the 1-Grassmanian bundle of the tangent bundle of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$, i.e. the fiber above a point $x\in\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$ consists of all complex lines in $T_{x}\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$. Then there is a natural biholomorphism $F:\operatorname{\mathcal{F}}_{1,2}(\operatorname{\mathbb{C}}^{d})\rightarrow\operatorname{\mathrm{Gr}}_{1}(T\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d}))$ given by
$F(x)=(x^{1},T_{x^{1}}\operatorname{\mathbb{P}}(x^{2})).$
Since $\Lambda_{1}(\Gamma)$ is a complex manifold,
$N:=\{(x,T_{x}\Lambda_{1}(\Gamma)):x\in\Lambda_{1}(\Gamma)\}$
is a complex submanifold of $\operatorname{\mathrm{Gr}}_{1}(T\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d}))$. By Equation (5),
$F^{-1}(N)=\Lambda_{1,2}(\Gamma)$
and so $\Lambda_{1,2}(\Gamma)$ is a complex submanifold. ∎
@@
We will now give a sketch of our original, more dynamical proof prior to the simplification suggested by Andres Sambarino. The argument requires that the Zariski closure of $\Gamma$ is simple.
@@
By Theorem 3.1, the $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$ is Ahlfors regular and the $\alpha_{1}$-Patterson–Sullivan measure of $\hat{\Gamma}$ on $\mathbb{CP}^{1}$ is mutually absolutely continuous with the Lebesgue measure. Since quasi-conformal maps preserve the Lebesgue measure class, we can then conclude that their pull-backs to $\partial\Gamma$ are mutually absolutely continuous. Then a result of Blayac–Canary–Zhu–Zimmer [5, Thm. 13.2] (see also Sambarino [29, Cor. 3.3.4]) implies that
@@
We then consider the identity component $\mathsf{G}^{0}$ of the real Zariski closure $\mathsf{G}$ of $\Gamma$. Assuming that $\mathsf{G}^{0}$ is simple, we may use an argument of Dal’bo–Kim [14] to show that $\mathsf{G}^{0}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$. Once we have done so, we can complete the proof of Theorem 4.1 just as above.
@@ Appendix A Reducing to the projective case
Notice that $E(x)^{1}$ and $E(x)^{2}$ only depend on $x^{k-1},x^{k},x^{k+1}$. Hence there exists a unique map $E_{k}:\operatorname{\mathcal{F}}_{k-1,k,k+1}(\operatorname{\mathbb{K}}^{d})\rightarrow\operatorname{\mathcal{F}}_{1,2}(\operatorname{\mathbb{K}}^{\bar{d}})$ satisfying
$E_{k}(x^{k-1},x^{k},x^{k+1})=(E(x)^{1},E(x)^{2})$
for all $x\in\operatorname{\mathcal{F}}_{\theta_{k}}(\operatorname{\mathbb{K}}^{d})$. Moreover, the formula for $E(x)^{1}$ and $E(x)^{2}$ imply that $E_{k}$ is a $\eta_{k}$-equivariant smooth embedding, which is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$.
The formula for $E(x)^{1}$ implies that $E_{k}$ descends to a smooth embedding $E_{k}^{1}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\rightarrow\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{k}})$. Hence (3) is true.
Next we prove (1). By (VII), the fact that $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-divergent implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-divergent, and $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))=E_{\theta_{k}}(\Lambda_{\theta_{k}}(\Gamma))$. In particular, $\Lambda_{1}(\eta(\Gamma))=E_{k}(\Lambda_{k}(\Gamma))$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-transverse, (III) implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-transverse. Since $E_{\theta_{k}}$ is an $\eta$-equivariant embedding, the fact that every point in $\Lambda_{\theta_{k}}(\Gamma)$ is a conical point for the $\Gamma$-action implies that every point in $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))$ is a conical point for the $\tau(\Gamma)$-action. Hence, $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-Anosov.